You conduct the following hypothesis test using the Spring Break 2020 data of GVSU students who went to Florida during spring break:

Ho: mu >= $675

Ha: mu < $675

The test statistic value is t = -3.31. What is the p-value for this test? Round to 3 decimal places.
Answer #1

Solution:-

n=20

df=n-1=20-1=19

Ho: μ $675

Ha: μ < $675

Test statistic is,

t = -3.31

P-value=P(t<t=-3.31)

P-value=0.002